OpenStudy (anonymous):

I am studying differentiation and integration on the circle and have difficulty with the slope which appears to simplify to the Cotangent function. By implication Integration on Cot(x) should indicate the circle as its area function however this is not working out - any assistance or clarification is greatly appreciated

OpenStudy (anonymous):

can you type out the question

OpenStudy (anonymous):

Based upon the circle equation x^2 + y^2 = 1 my attempt at dy/dx produced the following - x / ((1-x^2)^0.5). Alternatively if the same equation is presented in parametric form it would be x= cos(t) y = sin(t). If the latter is differentiated it produces -cos(t)/sin(t) as the slope which is -Cot(t). I am seeking to get a better appreciation of what happens should one wish to integrate the equation of the slope in either form ie. Integral of (-x / (1-x^2)^0.5 ) or Integral of -Cot(t) and whether this is consistent with the conventional outcome as regards derivation of Area. Hope this detail is feasible to follow. Regards Ed

OpenStudy (anonymous):

one second

OpenStudy (anonymous):

it would be nice to integrate a parametrically defined curve

OpenStudy (anonymous):

to integrate -x / sqrt (1-x^2) you use trig substitution I believe

OpenStudy (anonymous):

yeah thats easier to integrate the parametric integral

OpenStudy (anonymous):

cot x is easy to integrate let u = sin t, so we get -ln (sin(t)) , then plug in the limits

OpenStudy (anonymous):

so if you want to integrate from x= 0 to x = 1, we have 0 = cos(t) ---> t = pi/2, 1= cos (t) --> t = 0, so this will flip the integral

OpenStudy (anonymous):

ok i did it wrong

OpenStudy (anonymous):